OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. -1, 2, and 1 - i

OpenStudy (anonymous):

so, just double checking here, but the factors (aka zeros) of this polynomial are -1, 2 and (1-i), the last factor being a complex number?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

Hint: complex zeros always come in conjugate pairs (a+bi, a-bi)

OpenStudy (whpalmer4):

at least that's true if you want to have only real coefficients....

OpenStudy (anonymous):

Yeah i got that, i got to the point that i know it would be something like this, (x−1)(x+2)(x−(1+i))(x−(1−i). But i have no idea on how to actually write it in standard form.

OpenStudy (whpalmer4):

Multiply that sucker out! :-)

OpenStudy (whpalmer4):

Hint: do the two complex binomials first

OpenStudy (whpalmer4):

But before you do, remember that the zeros are the values of x that make one of the (x-zn) terms = 0. So if x = -1 is a zero, that means that the binomial is actually (x+1) so that (-1 + 1) = 0

OpenStudy (anonymous):

Well im not going to lie, i havent been paying any attention in math, and thats why im here haha. So i would appreciate a bit more guidance for this because i honestly have no idea how to multiply the two complex binomials and where to even start to do it.

OpenStudy (whpalmer4):

In other words, to construct your list of product terms, it's always (x - <zero>), not (x + <zero>)

OpenStudy (whpalmer4):

Okay, we can walk through it. But first, let's construct the right expression to multiply so that we don't have to do it over :-)

OpenStudy (whpalmer4):

So, just the real roots, x = -1, and x = 2, what would the polynomial look like (before multiplication)?

OpenStudy (anonymous):

Haha i have no idea. Im pretty much 100% lost in math right now, but im going to take a guess and say, (x−1)(x+2)?

OpenStudy (whpalmer4):

Nope, but good try :-) Remember, the reason they are called zeros is that the polynomial has a value of 0 at that point. Any polynomial can be written as \[(x - z_0)(x-z_1)(x-z_2)... = 0\] where \(z_0, z_1, z_2, ...\) are the zeros. Now, for the polynomial to equal 0, obviously one of those \(x-z_n)\) suckers has to equal 0, right?

OpenStudy (whpalmer4):

Oops, make that \((x-z_n)\)

OpenStudy (anonymous):

Yeah makes sense

OpenStudy (whpalmer4):

So if we a function that = 0 when x = -1, and when x = 2, that means there's an \((x+1)\) and a \((x-2)\) in there to cause that.

OpenStudy (anonymous):

Okay.

OpenStudy (whpalmer4):

So our polynomial having roots x=-1, x = 2, x = 1-i and x = 1+i is going to look like this before multiplication: \[(x+1)(x-2)(x-(1-i))(x-(1+i))\]

OpenStudy (anonymous):

Alright

OpenStudy (whpalmer4):

You know how to multiply \[(x+1)(x-2) = x^2 - 2x + x -2 = x^2 -x -2\] so that isn't a big deal. But how to multiply \[(x-(1-i))(x-(1+i))\]?

OpenStudy (anonymous):

Yeah i have no idea how to do that second part

OpenStudy (anonymous):

Hint: Don't get rid of the brackets inside.

OpenStudy (anonymous):

paranthesis*

OpenStudy (whpalmer4):

Well, same way, just somewhat more painful! \[(x-(1-i))(x-(1+i)) = x^2 - (1+i)x -(1-i)x -(-(1-i)(1+i))\] Now let's take a little side trip and evaluate \[(1+i)(1-i) = 1^2 -i + i - i^2 = 1-i^2\]but remember \(i^2 = -1\) so that becomes \[1 -(-1) = 2\] \[x^2 -(1+i)x - (1-i)x + 2\] \[x^2 - x -ix -x + ix + 2\] \[x^2 -2x + 2\]

OpenStudy (whpalmer4):

Now we multiply that with our two real zeros: \[(x^2 - 2x + 2)(x^2 - x -2)\]

OpenStudy (whpalmer4):

Give it a shot and let's see what you get!

OpenStudy (anonymous):

(1+i)(1−i)=12−i+i−i2=1−i2 I understand how you did that, and this, but remember i2=−1 so that becomes 1−(−1)=2, but i have no idea how you got this, x2−(1+i)x−(1−i)x+2. and everything afterwards

OpenStudy (whpalmer4):

Here, let's try it a slightly different way. Let's set c = 1+i and d = 1-i, then our expression is \[(x-c)(x-d) = x^2 -cx -dx + cd = x^2 -(c+d)x + cd\]Right?

OpenStudy (anonymous):

Oh so this is like a equation already setup? And we just plug the numbers in?

OpenStudy (whpalmer4):

But if c = 1+i and d = 1-i, then \[c+d = 1+i + 1 - i = 2\]and\[cd = (1+i)(1-i) = 1-i+i-i^2 = 2\]

OpenStudy (whpalmer4):

so that makes the product of \[(x-c)(x-d) = x^2 - (2)x + 2\] Much easier that way!

OpenStudy (whpalmer4):

What remains is to multiply\[(x^2-2x+2)(x^2-x-2)\]

OpenStudy (whpalmer4):

\[x^4-x^3-2x^2-2x^3+2x^2+4x+2x^2-2x-4 = x^4-3x^3+2x^2+2x-4\]

OpenStudy (anonymous):

Yeah that last step is easy

OpenStudy (whpalmer4):

Now the fun begins :-) We ought to check our work! x=2 \[2^4-3(2^3)+2(2)^2+2(2)-4 = 16 - 24 + 8 + 4 - 4 = 0 \]That one works! x=-1 \[(-1)^4-3(-1)^3+2(-1)^2 + 2(-1)-4 = 1 -(-3) + 2(1) - 2 -4 = 0 \]That one works!

OpenStudy (whpalmer4):

The other two are a bit more tedious: \[(1-i)^2=(1 -2i + i^2) = -2i\]\[(1-i)^3 = -2i(1-i) = -2i + 2i^2 = -2i -2\]\[(1-i)^4 = ((1-i)^2)^2 = (-2i)^2 = 4i^2 = -4\] \[(1-i)^4 - 3(1-i)^3 + 2(1-i)^2 + 2(1-i) - 4 = \]\[-4 -3(-2i-2) + 2(-2i) + 2(1-i) - 4 = -4 +6i + 6 -4i + 2 - 2i - 4 = \]\[-4+6+2-4+6i-4i+2i =0\] I'll let you do the other one as a learning exercise :-)

OpenStudy (anonymous):

.This is soo much for one problem...

OpenStudy (whpalmer4):

@#$#@% I typed the wrong sign in front of that last 2i in the final line

OpenStudy (whpalmer4):

It's good for you, it grows hair on your chest as my old math teacher used to say :-)

OpenStudy (anonymous):

Haha, well im going to take a break. I feel like just getting this answer without knowing anything of nothing, is good enough for now. I appreciate your help a lot! Thanks!

OpenStudy (whpalmer4):

Sure, I'm ready for a break too! Shoot message any time if you need some help. I'm not always online, but if you're interested in learning, I'm interested in helping.

OpenStudy (whpalmer4):

"shoot me a message"

OpenStudy (anonymous):

Will do! :)